I am trying to find out how to solve this redacted problem. What is the process for finding the answers to this question? Even if realistic numbers are used in the steps in the solution?

1. A steel wire is connected at joints A and B and looped over a lightweight pulley as shown below. It is connected in a way in which it has no slack or tension at a temperature of 23° C. The steel wire has a modulus of elasticity ofGPa, yield strength of MPa, and an ultimate strength of 420 MPa. The wire also has a diameter of 4 mm and a coefficient of thermal expansion of 1.1x105 c1. Joint B can be moved downwards by rotating a nut on a square threaded screw with a lead of mm. A. Joint B is moved lower by two full turns of the nut. What is the stress inside the wire? The threads of the screw are prevented from turning and the temperature remains at 23°0C B. After stretching the wire according to part A, the temperature drops to·rc. Determine force the wire exerts on Joint A at this new temperature. 0.2 m immovable nut threaded bolt
Answer #1

In this question some of the data are blackened and thus we have to assume suitably to solve the problem. If you want to know the steps about how to proceed solving these kind of problems. Then, steps are as follows:

1) Draw a free body diagram as shown in figure 1.

belt

2) As the nut is screwed, the steel wire get stressed and the stress is found by the formula:

Stress = \frac{Force}{Area} = \frac{Tension in steel wire}{cross sectional area of wire}

Stress = T/0.785*d^2 (Answer A)

Where, d is the diameter of the steel wire.

Anwer B)

As the temperature decreases, the tension in the wire increases and thus the force exerted on joint A will also increase.